OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x − 3y − z if x = −2, y = 3, and z = −2 −11 11 −15 15

OpenStudy (anonymous):

just plug it in, 2(-2) - 3(3) -(-2) = -4-9+2 = -11

OpenStudy (anonymous):

plug in the numbers 2(-2) - 3(3) - (-2) -4 - 9 + 2 = -11

OpenStudy (anonymous):

Subsitution... 2(-1) - 3(3) - (-2) The use of ( ) whenever you substitute is important and a good practice to get into the habit of doing now. :-)

OpenStudy (anonymous):

-11 is answer